# Hyperbolic Random Graphs: Clique Number and Degeneracy with Implications for Colouring

### v1

@@ 4.1 The gap between Clique Number and Degeneracy
Case 3 [$r\in R/2+\Theta(1)$]: Using Lemmas 4.1 and 3.2, we get w.e.h.p. that $|K|\leq(1-\varepsilon)|\operatorname{\Gamma}(u)|\leq(1-\varepsilon)|\operatorname{\Gamma}(u^{*})|\leq(1-\varepsilon)\kappa(G)$.
@@
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then there exists a constant $\varepsilon\in(0,1)$ such that with high probability, $\omega(G)\leq(4(1-\varepsilon)/3)^{\alpha}\sigma(G).$
@@ 4.2 Cliques larger than the Core
In this section, we show that there exist a super-constant number of unique cliques that contain the core, but are strictly larger than it. The overall argument goes as follows. We consider a set of vertices with radial coordinates slightly outside the core. We show that any vertex of this set has a constant probability to be adjacent to all vertices belonging to the core, and thus induces a clique that is larger than the core itself.
@@
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG Then w.e.h.p. there exist $\omega(\log(n))$ cliques that are larger than $\sigma(G)$.
Proof 4.8.
Figure 4: Sketch of the proof idea for Proposition 4.7. (a) Illustration of the set of points $\mathcal{A}$ (blue) of width $\xi$ slightly outside of the core (pink). The intersection with a sector of angle $\psi$ and the band $\mathcal{A}$ forms a box $\mathcal{A}_{i}$ that contains a vertex w.e.h.p. The number of non-intersecting boxes is $k=2\pi/\psi\in\omega(\log(n))$. (b) A box $\mathcal{A}_{i}$ that contains vertex $u$. The hatched area $\mathcal{F}_{u}$ with angle at most $\phi$ is the corresponding forbidden area of $u$. Any point in $\operatorname{\mathcal{F}_{u}}$ has distance at least $R$ to $u$. An adequate choice for the width $\xi$ yields that this area is empty with constant probability.
@@
$\begin{split}\mu(\mathcal{A})&=\int_{R/2}^{R/2+\xi}\frac{\sinh(\alpha x)}{\cosh(\alpha R)-1}=\frac{\cosh(\alpha(R/2+\xi))-\cosh(\alpha R/2)}{\cosh(\alpha R)-1}\\ &=(1+\Theta(e^{-\alpha R}))(e^{-\alpha(R/2-\xi)}-e^{-\alpha R/2})\\ &=(1+\Theta(e^{-\alpha R}))(n^{-\alpha}e^{-\alpha C/2}{\left(1+\log^{3}(n)n^{\alpha-1}\right)}-n^{-\alpha}e^{-\alpha C/2})\\ &=(1+\Theta(e^{-\alpha R}))\big{(}e^{-\alpha C/2}\log^{3}(n)n^{-1}\big{)}\in\Theta\left(\log\left(n^{3}/n\right)\right).\end{split}$ (3)
@@
In the second step, we show that for any vertex $u\in A$, the probability that there exists a vertex in its forbidden area $\operatorname{\mathcal{F}_{u}}$ is strictly less than $1$. Since $r(u)\leq R/2+\xi$, we have that $u\in A$ has an edge to any vertex in the area $\mathcal{B}_{0}(R/2-\xi)$. Hence, we have $\operatorname{\mathcal{F}_{u}}\subset\mathcal{B}_{0}(R/2)\setminus\mathcal{B}_{0}(R/2-\xi)$. Now, for $R/2\geq x\geq R/2-\xi$ we seek to find the angle size $\phi=2(\pi-\theta_{R}(r(u),x))$ in order to upper bound $\mu(\operatorname{\mathcal{F}_{u}})$. Since $\theta_{R}(r,x)$ is monotonically decreasing in $x$ we have $\phi\leq 2(\pi-\theta_{R}(R/2+\xi,R/2))$ and we obtain
@@
where we used $\int_{R/2-\xi}^{R/2}\rho(x)\,\mathrm{d}x\leq\int_{R/2}^{R/2+\xi}\rho(x)\,\mathrm{d}x=\mu(\mathcal{A})$. By our choice of $\xi=\log(1+\log^{3}(n)n^{\alpha-1})\in o(1)$, we can apply Lemma B.5 and obtain $\theta_{R}(R/2+\xi,R/2)\geq\pi-2\sqrt{\frac{\log^{3}(n)}{n^{1-\alpha}}}$. In Equation 3 we established that $\mu(\mathcal{A})\in\Theta\left(\log^{3}\left(n\right)/n\right)$. Combining the two leads to $\mu(\operatorname{\mathcal{F}_{u}})\in\mathcal{O}\Big{(}\sqrt{\log^{9}(n)n^{\alpha-3}}\Big{)}$.
@@
We now establish our third and final desired property. Namely, we construct a subset $U\subset A$, consisting of vertices whose forbidden areas are disjoint and for which $\mathbb{E}\left[|U|\right]\in\omega(\log(n))$.
@@
To wrap things up, recall that in the first step we established that each sector $\mathcal{A}_{i}$ contains a vertex w.e.h.p. Moreover, since the $X_{i}$ are independent, we have by a Chernoff bound (Theorem A.1) that $X\in\omega(\log(n))$ w.e.h.p. Though these two events are not independent, we can apply the union bound to their complements to obtain that w.e.h.p. $\omega(\log(n))$ vertices outside of $\mathcal{B}_{0}(R/2)$ are adjacent to all vertices in $\mathcal{B}_{0}(R/2)$, which finishes the proof.
@@ 4.3 Upper Bound on the Clique Number
Before going into details, we lay out our proof strategy. We aim to bound the region where a clique can be located. Since vertices are adjacent if and only if their (hyperbolic) distance is at most $R$, this can be done by characterising a shape that covers any hyperbolic region of diameter $R$. A classic result by Jung [19] answers the question of how large the radius of a ball in Euclidean space needs to be at most, so that its interior can contain an entire set of points of fixed diameter. The hyperbolic version of this result was discovered by Dekster [13, 14] nearly a century later. He extended Jung’s result to (among other geometries) hyperbolic space and we apply it as follows: we identify $\mathcal{O}(n^{3})$ many balls where one of these balls contains the clique of largest of size $\omega(G)$. This clique (and all the other identified cliques) needs to be located in a ball $B_{x}(r)$ with radius $r$ large enough. We use the hyperbolic variant of Jung’s theorem to upper bound $r$ which, in turn, allows us to upper bound the area of this ball. This yields an upper bound for the amount of vertices one such ball $B_{x}(r)$ could contain w.e.h.p., leading to an upper bound for $\omega(G)$. Since we need only consider at most $\mathcal{O}(n^{3})$ balls, a union bound is sufficient to derive the same bound for the worst case. We work with the following version of Jung’s theorem for hyperbolic geometry.
@@
for arbitrary $c$. Thus to finish the proof we need to show concentration, which via a union bound over all triplets $u,v,w$ will yield the result. To show concentration we apply a Chernoff bound. Using $\varepsilon=(1/\log(n))$ we obtain
$\mathbb{P}\left(Y_{\mathcal{B}}>(1+1/\log(n))\mathbb{E}[Y_{\mathcal{B}}]\right)$ $\leq e^{-(1/\log(n))^{2}/3\cdot\mathbb{E}[Y_{\mathcal{B}}]}\leq e^{-\Theta(1)(n^{1-\alpha})/\log^{2}(n)}\leq n^{-c}$
@@ 5 Geometric Inhomogeneous Random Graphs
Because of this relationship, GIRGs are used as proxies for HRGs in some theoretical and experimental works. This is partly done because GIRGs are (by design) far more tractable than HRGs. It is therefore valuable to understand differences between the two models. In [5] experimental evidence was given to suggest that the “sandwiching” of an HRG by two standard GIRGs is not tight. In Corollary 5.5 we provide theoretical result demonstrating a difference between the two models.
@@
    Every $u\in V$ is equipped with a random tuple $(w_{u},x_{u})$, where $w_{u}\in[1,\infty)$ has density $f(y)=(\beta-1)y^{-\beta}$ and $x_{u}$ is drawn uniformly at random from $[0,1]$;
@@
One way of thinking of a GIRG is that vertices are being thrown uniformly at random onto the 1-dimensional torus $\mathbb{T}^{1}$, and connected according to whether their distance is below threshold $t(u,v)$. Analogously to HRGs, for a vertex $u$ of a GIRG we define the inner-degree of $u$ to be $|\operatorname{\Gamma}(u)|=|\{v\in V|\text{$u$ and $v$ are connected and }w_{v}\geq w_{u}\}|.$ The proofs of Sections 3 and 3.2 can be adapted to the GIRG model to characterize the degeneracy via the largest inner-degree.
@@
Proof 5.4 (Proof sketch).
We prove the statement in three steps. First, we compute the probability for a vertex $v$ to lie in the inner-neighbourhood of a vertex $u$ depending on the weight $w_{u}$. We then use this to find the weight $w^{*}$ maximising expected inner-neighbourhood. Combining the two, we use $w^{*}$ to obtain upper- and lower bounds on the largest inner-degree of a vertex w.e.h.p. The statement then follows from Corollary 5.2.
@@
Corollary 5.5.
Fix an $\alpha\in(1/2,1)$, and let $G\sim\mathcal{G}(n,2\alpha+1,\lambda)$ be a standard GIRG and $H\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then w.e.h.p.
@@ 6 Conclusion
  • Is the chromatic number bounded away from the degeneracy, the clique number, or from both?
@@
  • Can the gap between upper and lower bound for the clique number be closed?
There are several further directions of research such as determining further differences between HRGs and GIRGs or designing colouring and other algorithms for HRGs in various models of computation.
Appendix A Concentration bounds
We frequently use the following Chernoff bounds [30, Theorem 4.4].
Theorem A.1 (Chernoff bound).
$\mathbb{P}\left(X\geq(1+\varepsilon)\mathbb{E}[X]\right)\leq e^{-\varepsilon^{2}/3\cdot\mathbb{E}[X]}\text{ and }\mathbb{P}\left(X\leq(1-\varepsilon)\mathbb{E}[X]\right)\leq e^{-\varepsilon^{2}/2\cdot\mathbb{E}[X]}.$
Appendix B Bounds on $\theta_{R}(\cdot,\cdot)$
Before we give bounds on the measure of areas in Appendix C, we provide some refined bounds on $\theta_{R}(r,q)$ for several pair of radii with $r+q\geq R+\Delta$ for specific $\Delta\in\Theta(1)$. This will prove useful since the considered inner-neighbourhoods belong to vertices that are close to the core ball $\mathcal{B}_{0}(R/2)$. We use the following inequality for the $\arccos$.
Lemma B.1.

### v2

@@ 4.1 The gap between Clique Number and Degeneracy
Case 3 [$r_{K}\in R/2+\Theta(1)$]: Let $\xi\in o(1)$. Using Lemmas 4.1 and 3.2, we get w.e.h.p. that
$\omega(G)=|K|\leq\frac{1-\varepsilon}{1-\xi}|\operatorname{\Gamma}(u)|\leq\frac{1-\varepsilon}{1-\xi}|\operatorname{\Gamma}(u^{*})|\leq(1-\varepsilon)\kappa(G),$
for an adequate choice of $\xi$.
@@
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then there exists a constant $\varepsilon\in(0,1)$ such that w.e.h.p., $\omega(G)\leq(4(1-\varepsilon)/3)^{\alpha}\sigma(G).$
@@ 4.2 Cliques larger than the Core
In this section, we show that there exist a super-constant number of unique cliques that contain the core, but are strictly larger than it. The overall argument goes as follows. We consider a set of vertices with radial coordinates slightly outside the core. We show that any vertex of this set has a constant probability to be adjacent to all vertices belonging to the core, and thus induces a clique that is larger than the core itself. To this end, the following lemma concerned about points close to the core proves useful.
Lemma 4.7.
Thus, when $r$ approaches $R/2$ from above, the angle $\theta_{R}(r,R/2)$ approaches $\pi$ from below.
@@
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then w.e.h.p. there exist $\omega(\log(n))$ cliques that are larger than $\sigma(G)$.
Proof 4.9.
Figure 4: Sketch of the proof idea for Proposition 4.8. (a) Illustration of the set of points $\mathcal{A}$ (blue) of width $\xi$ slightly outside of the core (pink). The intersection with a sector of angle $\psi$ and the band $\mathcal{A}$ forms a box $\mathcal{A}_{i}$ that contains a vertex w.e.h.p. The number of non-intersecting boxes is $k=2\pi/\psi\in\omega(\log(n))$. (b) A vertex $u$ located on the boundary of the area $\mathcal{A}$. The hatched area $\mathcal{F}_{u}$ with angle at most $\phi$ is the corresponding forbidden area of $u$. Any point in $\operatorname{\mathcal{F}_{u}}$ has distance at least $R$ to $u$. An adequate choice for the width $\xi$ yields that this area is empty with constant probability.
@@
$\begin{split}\mu(\mathcal{A})&=\int_{R/2}^{R/2+\xi}\frac{\sinh(\alpha x)}{\cosh(\alpha R)-1}=\frac{\cosh(\alpha(R/2+\xi))-\cosh(\alpha R/2)}{\cosh(\alpha R)-1}\\ &=(1+\Theta(e^{-\alpha R}))(e^{-\alpha(R/2-\xi)}-e^{-\alpha R/2})\\ &=(1+\Theta(e^{-\alpha R}))(n^{-\alpha}e^{-\alpha C/2}{\left(1+\log^{3}(n)n^{\alpha-1}\right)}-n^{-\alpha}e^{-\alpha C/2})\\ &=(1+\Theta(e^{-\alpha R}))\big{(}e^{-\alpha C/2}\log^{3}(n)n^{-1}\big{)}\in\Theta\left(\log^{3}\left(n\right)/n\right).\end{split}$ (3)
@@
In the second step, we show that for any vertex $u\in A$, the probability that there exists a vertex in its forbidden area $\operatorname{\mathcal{F}_{u}}:=\{x\in\mathcal{B}_{0}(R/2):\operatorname{d_{h}}(u,x)>R\}$ (see Figure 4 b) is strictly less than $1$. Since $r(u)\leq R/2+\xi$, we have that $u\in A$ has distance at most $R$ (and thus, an edge) to any vertex in the area $\mathcal{B}_{0}(R/2-\xi)$. Hence, we have $\operatorname{\mathcal{F}_{u}}\subset\mathcal{B}_{0}(R/2)\setminus\mathcal{B}_{0}(R/2-\xi)$. Now, for $R/2\geq x\geq R/2-\xi$ we seek to find the angle size $\phi=2(\pi-\theta_{R}(r(u),x))$ in order to upper bound $\mu(\operatorname{\mathcal{F}_{u}})$. Since $\theta_{R}(r,x)$ is monotonically decreasing in $x$ we have $\phi\leq 2(\pi-\theta_{R}(R/2+\xi,R/2))$ and we obtain
@@
where we used $\int_{R/2-\xi}^{R/2}\rho(x)\,\mathrm{d}x\leq\int_{R/2}^{R/2+\xi}\rho(x)\,\mathrm{d}x=\mu(\mathcal{A})$. By our choice of $\xi=\log(1+\log^{3}(n)n^{\alpha-1})\in o(1)$, we can apply Lemma 4.7 and obtain $\theta_{R}(R/2+\xi,R/2)\geq\pi-2\sqrt{\frac{\log^{3}(n)}{n^{1-\alpha}}}$. In Equation 3 we established that $\mu(\mathcal{A})\in\Theta\left(\log^{3}\left(n\right)/n\right)$. Combining the two leads to $\mu(\operatorname{\mathcal{F}_{u}})\in\mathcal{O}\Big{(}\sqrt{\log^{9}(n)n^{\alpha-3}}\Big{)}$.
@@
We now establish our third and final desired property. Namely, we construct a subset $U\subset A$, consisting of vertices whose forbidden areas are empty and disjoint and for which $\mathbb{E}\left[|U|\right]\in\omega(\log(n))$.
@@
To wrap things up, recall that in the first step we established that each sector $\mathcal{A}_{i}$ contains a vertex w.e.h.p. Moreover, since the $X_{i}$ are independent, we have by a Chernoff bound that $X\in\omega(\log(n))$ w.e.h.p. Though these two events are not independent, we can apply the union bound to their complements to obtain that w.e.h.p. $\omega(\log(n))$ vertices outside of $\mathcal{B}_{0}(R/2)$ are adjacent to all vertices in $\mathcal{B}_{0}(R/2)$, which finishes the proof.
@@ 4.3 Upper Bound on the Clique Number
Before going into details, we lay out our proof strategy. We aim to bound the region where a clique can be located. Since vertices are adjacent if and only if their (hyperbolic) distance is at most $R$, this can be done by characterising a shape that covers any hyperbolic region of diameter $R$. A classic result by Jung [19] answers the question of how large the radius of a ball in Euclidean space needs to be at most, so that its interior can contain an entire set of points of fixed diameter. The hyperbolic version of this result was discovered by Dekster [13, 14] nearly a century later. He extended Jung’s result to (among other geometries) hyperbolic space and we apply it as follows: we identify $\mathcal{O}(n^{3})$ many balls where one of these balls contains the clique of largest size $\omega(G)$. This clique (and all the other identified cliques) needs to be located in a ball $B_{x}(r)$ with radius $r$ large enough. We use the hyperbolic variant of Jung’s theorem to upper bound $r$ which, in turn, allows us to upper bound the area of this ball. This yields an upper bound for the amount of vertices one such ball $B_{x}(r)$ could contain w.e.h.p., leading to an upper bound for $\omega(G)$. Since we only need to consider at most $\mathcal{O}(n^{3})$ balls, a union bound is sufficient to derive the same bound for the worst case. We work with the following version of Jung’s theorem for hyperbolic geometry.
Theorem 4.10.
$D\geq 2\sinh^{-1}\left(\sqrt{\frac{d+1}{2d}}\sinh(r)\right).$
@@
Proof 4.12.
Using that $d=2$, we directly get from Theorem 4.10 for diameter $R$ that $\frac{R}{2}\geq\sinh^{-1}(\sqrt{3/4}\sinh(r)).$ Rearranging and using for $x\in\mathbb{R}$ that $\sinh(x)=\frac{1}{2}e^{x}(1-e^{-2x})$ yields
@@
for arbitrary $c$. Thus to finish the proof we need to show concentration, which via a union bound over all triplets $u,v,w$ will yield the result. To show concentration we apply a Chernoff bound Theorem 2.1. Using $\varepsilon=(1/\log(n))$ we obtain
$\mathbb{P}\left(Y_{\mathcal{B}}>(1+1/\log(n))\mathbb{E}[Y_{\mathcal{B}}]\right)$ $\leq e^{-\mathbb{E}[Y_{\mathcal{B}}]/(3(1/\log(n))^{2})}\leq e^{-\Theta(1)(n^{1-\alpha})/\log^{2}(n)}\leq n^{-c}$
@@ 5 Geometric Inhomogeneous Random Graphs
Because of this relationship, GIRGs are used as proxies for HRGs in some theoretical and experimental works. This is partly done because GIRGs are (by design) far more tractable than HRGs. It is therefore valuable to understand differences between the two models. In [5] experimental evidence was given to suggest that the “sandwiching” of an HRG by two standard GIRGs is not tight. In Corollary 5.5 we provide a theoretical result demonstrating a difference between the two models.
@@
    Every $u\in V$ is equipped with a random tuple $(w_{u},x_{u})$, where weight $w_{u}\in[1,\infty)$ has density $f(y)=(\beta-1)y^{-\beta}$ and coordinate $x_{u}$ is drawn uniformly at random from $[0,1]$;
@@
One way of thinking of a GIRG is that vertices are being thrown uniformly at random onto the 1-dimensional torus $\mathbb{T}^{1}$, and connected according to whether their distance is below their threshold $t(u,v)$. The weights are drawn according to a Pareto distribution. Analogously to HRGs, for a vertex $u$ of a GIRG we define the inner-degree of $u$ to be $|\operatorname{\Gamma}(u)|=|\{v\in V|\text{$u$ and $v$ are connected and }w_{v}\geq w_{u}\}|.$ The proofs of Sections 3 and 3.2 can be adapted to the GIRG model to characterise the degeneracy via the largest inner-degree.
@@
Proof 5.4.
We bound the maximal inner-degree $|\operatorname{\Gamma}(u^{*})|$; the statement then follows from Corollary 5.2. Notice that, independent of the geometric distance, a vertex with weight $w$ is adjacent to any vertex with weight $w^{\prime}$ if $w^{\prime}\geq\frac{n}{w\lambda}$ since $t(w,w^{\prime})=\frac{\lambda ww^{\prime}}{2n}\geq\frac{\lambda w\frac{n}{w\lambda}}{2n}=\frac{1}{2}$ which is the maximal distance between two points in the unit torus. Thus, using $\beta=2\alpha+1$, the probability that a vertex $v$ is in the inner-neighbourhood of a vertex $u$ with weight $w$ is
$\mathbb{P}\left(v\in\Gamma(u)\right)$ =($=\mathbb{P}\left(\{\{u,v\}\in E\}\cap\{W_{v}\geq w\}\right)$
[…]
Next, we calculate the value $w^{*}$, which maximises the expected inner-degree. To this end, we consider the probability measure of the inner-neighbourhood, take its derivative with respect to $w$ and set it equal to 00. Differentiating yields
$\frac{\mathrm{d}}{\mathrm{d}w}\mathbb{P}\left(v\in\Gamma(u)\right)$ $=\frac{2(nw)^{-(2\alpha+1)}(1-\alpha)\alpha\left(n^{2\alpha}w^{2}\lambda-nw^{4\alpha}\lambda^{2\alpha}\right)}{2\alpha-1},$
and solving for the maximum reveals $w^{*}=(2(1-\alpha))^{\frac{1}{4\alpha-2}}\sqrt{\frac{n}{\lambda}}\in\Theta(\sqrt{n}).$
We plug in $w^{*}$ for the weight of $u$ denoted by $u^{*}$ into $\mathbb{P}\left(v\cap\Gamma(u^{*})\right)$ and get by (5.4) that
Recalling that $\sigma(G)=(1\pm o(1))\lambda^{-\alpha}n^{1-\alpha}$ w.e.h.p., the upper bound now follows from the expectation of $|\operatorname{\Gamma}(u^{*})|$ and applying a Chernoff bound in conjunction with a union bound. The lower bound is established by showing that there exists a vertex within the range of weights $\tilde{w}=[w^{*}(1+n^{\alpha-1}\log^{2}(n))^{-1/(2\alpha)},w^{*}]$ w.e.h.p. and lower bound the inner-degree of such vertex. Using the Pareto distribution and $w^{*}\in\Theta(\sqrt{n})$, we calculate the probability for a vertex $u$ to belong to the range of weights $\tilde{w}$. We obtain
$\mathbb{P}\left(W_{u}\in\tilde{w}\right)$ $=\mathbb{P}\left(w^{*}(1+n^{\alpha-1}\log^{2}(n))^{-1/(2\alpha)}\leq W_{u}\leq w^{*}\right)$ (Range of $(\text{Range of }\tilde{w})$
[…]
$=(w^{*})^{-2\alpha}n^{\alpha-1}\log^{2}(n)$
$=\Theta(1)\frac{\log^{2}(n)}{n}.$ $(w^{*}\in\Theta(\sqrt{n}))$
By this we have $\mathbb{E}\left[|V\cap\tilde{w}|\right]\in\omega(\log(n))$. Using a Chernoff bound there exists a vertex within the desired weight range $\tilde{w}$ w.e.h.p. To conclude the proof we lower bound the inner-degree of a vertex $\tilde{u}$ included in the weight range $\tilde{w}$. Note that $\tilde{w}=(1-o(1))w^{*}=(2-o(1)(1-\alpha))^{\frac{1}{4\alpha-2}}\sqrt{\frac{n}{\lambda}}$. We then have via 5.4
@@
Corollary 5.5 (GIRG-HRG degeneracy difference).
Fix an $\alpha\in(1/2,1)$. Let $G\sim\mathcal{G}(n,2\alpha+1,\lambda)$ be a standard GIRG and $H\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then w.e.h.p.
@@ 6 Conclusion
  • Is the chromatic number bounded away from the degeneracy, the clique number, or both?
@@
There are further directions of research such as determining other differences between HRGs and GIRGs or designing colouring algorithms for HRGs in various models of computation.
Appendix A Appendix
A.1 Bounds on $\theta_{R}(\cdot,\cdot)$
Before we give bounds on the measure for of the area inner-ball Lemma 3.4 in Section A.2, we provide some refined bounds on $\theta_{R}(r,q)$ for several pair of radii with $r+q\geq R+\Delta$ for specific $\Delta\in\Theta(1)$. We use the following inequality for the $\arccos$.
Lemma A.1.